OpenStudy (anonymous):

evaluate the function at the given values of the variable. f(x)=the sq root of x-1 x=5,0

OpenStudy (anonymous):

\[\sqrt{5-1}\] = \[\sqrt{4}\] = 2 ---- \[\sqrt{0-1}\] = \[\sqrt{-1}\] = i

OpenStudy (anonymous):

what happened to x=5,0 ??

OpenStudy (anonymous):

Looks like you're evaluating two different values of x, so for the first you sub. 5 for x and the second you sub 0.